OpenStudy (anonymous):

Which of the following methods would be the easiest to use to solve x^2 – 6 = 0?

OpenStudy (anonymous):

What are the following methods to choose from?

OpenStudy (anonymous):

I don't see any given methods so i suggest you to do this equation normally. x^2 = 6\[x = \pm \sqrt{6}\]

OpenStudy (anonymous):

Which of the following methods would be the easiest to use to solve x2 – 6 = 0? A. factoring B. isolating the x2 term and finding the square root of both sides C. using the quadratic formula D. all three methods would be easy and effective

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

thank ya